The number of bacteria in a yeast culture grows at a rate which is proportional to the number present. If


the population of a colony of yeast bacteria triples in one hour, find the number of bacteria which will be


present at the end of five hours.

1
Expert's answer
2021-11-19T07:02:45-0500

"\displaystyle
\text{Let N be the no of bacteria and $\frac{dN}{dt}$ = rate of increase of bacteria}\\
\frac{dN}{dt}=kN\\
\implies \frac{dN}{N}= kdt\\
\ln N = kt + A\\
e^{\ln N} = e^{kt + A}\\
\implies N = ce^{kt}\\
\text{Let N = 3c, t = 1, then we have that}\\
\implies 3c = ce^{kt}\\
\therefore e^k =3\implies k = 1.0986\\
\text{Next, we find the no of bacteria present after 5 hours}\\
N = ce^{5k}\\
\implies nc = ce^{1.0986\cdot 5}\\
\implies n = e^{1.0986\cdot 5}= 243\\
\text{Therefore, the number of bacteria present after 5 hours is 243 times more than the }\\
\text{initial population of bacteria}"